average score is 42 and the standard deviation of scores is 24. Assume that the scores are




normally distributed answer the following questions




What is the probability that the candidates score




A. Exceed 65




B Between 40 and 60





Find the number of candidates whose score




A Exceed 407






1
Expert's answer
2022-05-04T14:54:24-0400

A.


"P(X>65)=1-P(Z\le\dfrac{65-42}{24})\approx0.1689"

B.


"P(40<X<60)=P(Z<\dfrac{60-42}{24})"

"-P(Z\le\dfrac{40-42}{24})\approx0.3066"

C.


"P(X>407)=1-P(Z\le\dfrac{407-42}{24})\approx0"

There are no candidates whose score exceed 407.